[Ian, Torrey]
We found the beam in transmission of the cavity when it flashes. Alignment of the beam into the cavity has so far been unsuccessful. This screenshot is an example of the beam we see. So far the alignment process has been:
1) Scan the cavity at large amplitude with the current to locate the REFL dip.
2) Lower the scan amplitude and adjust the temperature according such that the REFL dip stays in the scan.
3) Repeat until a scan is no longer needed and manual park the cavity on resonance using the temperature.
4) Attempt to adjust the input mirrors while parked on resonance.
There a few problems with this. First, the smallest step size we can take on the temperature is too large. Makes it difficult to park on resonance. Second the stability of the laser's frequency isn't the best for this process. When attempted to manually park on resonance it will often drift out of range in a few seconds. The next logical step would be to lock the cavity on this mode, but we have been unsuccessful in doing so. Either the loop shape isn't quite right or the mode resonant in the cavity is just too garbage to lock on.
Few more examples of things we saw: Aliasing effects, some garbage but was fixed due to aliasing effects of our scan frequency and the scanning slit frequency. Unsure what this is, we've seen it a couple of times but I'm not sure it is significant. Would be good to find out what kind of lenses are in the cavity chamber.